May has Php 7.50 in 10 cents and 25 cents coins.The number of 25 cents coins is n.If the number of 10 cents coins is 3 less than 4 times the number of 25 cents coins,how many coins of each denomination does she have?

Answer by addingup(3677)

0.25n+0.10(4n-3) = 7.50
0.25n+0.40n-0.30 = 7.50
0.65n = 7.80
n = 12 He has 12 0.25 coins = Php 3
The difference is 0.10 coins:
7.50-3 = 4.50; 4.50/0.10 = 45
He has 12 0.25 coins and 45 0.10 coins